OpenStudy (anonymous):

Evelyn purchased a dining room set with a 12-month deferred payment plan. The interest rate on the plan was 25.57%. The dining room set must be paid off within two years after the deferment period. What is Evelyn’s monthly payment after the deferment period if the price of the dining room set was $2,875 and she made no payments during the deferment period? $119.79 $154.28 $170.60 $198.68

OpenStudy (anonymous):

The interest rate on the plan was 25.57% so she must pay back 125.57% of the original price - $2,875

OpenStudy (anonymous):

total payments necessary are: $2,875 * 1.2557 , the small number is the %/100

OpenStudy (anonymous):

whats the formula? or is that it?

OpenStudy (anonymous):

the formula for principal+interest is: return = (1 + i )*principal , i is the interest rate hundredths

OpenStudy (anonymous):

can you help me figure out the answer please

OpenStudy (anonymous):

we 1) find the total repayment debt 2) find monthly payment

OpenStudy (anonymous):

for the formula they have interest rate in "hundredths" 10% corresponds to 10/100 i=0.1 25% corresponds to 25/100 i=0.25 100% corresponds to 100/100 i=1 we need this interest for 25.57%

OpenStudy (anonymous):

so .25?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

now formula was \[return = (1+i) \times principal\] the "principal" in this case is the cost of the set. The store invests it into Evelyn's household, so other direction - mechanism is the same.

OpenStudy (anonymous):

r=(1+.25) x 2875 ?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

3593.75

OpenStudy (anonymous):

that's the total money Evelyn must return to the store/credit institution

OpenStudy (anonymous):

"The dining room set must be paid off within two years" that's how many months with payments?

OpenStudy (anonymous):

24 months

OpenStudy (anonymous):

@mathessentials

OpenStudy (anonymous):

correct, so she got 24 months to repay the $3593.75

OpenStudy (anonymous):

than what

OpenStudy (anonymous):

they will set the minimum so every 24 months pay the same

OpenStudy (anonymous):

monthly payment = debt / months monthly payment = $3593.75 / 24

OpenStudy (anonymous):

149.74

OpenStudy (anonymous):

it's $150 when you take the precise 25.57% instead of just ~%25

OpenStudy (anonymous):

than what? or is that it?

OpenStudy (anonymous):

I'm not sure why our result is off by ~$4.5

OpenStudy (anonymous):

idk :( but is the answer 154.28

OpenStudy (anonymous):

I believe so, it's the closest one

OpenStudy (anonymous):

thank you! do you think you could try and help with more

OpenStudy (anonymous):

time for one more

OpenStudy (anonymous):

Zoey opened a savings account four years ago with a deposit of $935.15. The interest on the account compounds monthly. If Zoey’s current account balance is $1,128.16, what is the interest rate on the account? 0.4% 4.7% 14.2% 18.9%

OpenStudy (anonymous):

there's a formula for compounding: \[A=P(1+\frac{ i }{ n })^{ny}\] A: Account balance P: Principal i: interest y: number of years n: times it is compounded p.a.

OpenStudy (anonymous):

basically we just insert everything where it goes. data given was: A: $1,128 P: $935 y: 4 years n: 12 times p.a. (monthly)

OpenStudy (anonymous):

if it is compounded monthly, how often is it compounded in a year?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

yep, so n=12

OpenStudy (anonymous):

\[$1128=$935(1+\frac{ i }{ 12 })^{12 \times 4}\]

OpenStudy (anonymous):

4.7%

OpenStudy (anonymous):

Wyatt purchased a living room set for $4,109 with an 18-month deferred payment plan. The interest rate on the plan was 27.41%. The living room set must be paid off within three years after the deferment period. What is the total cost of the living room set if she made no payments during the deferment period and made only minimum payments after the deferment period? $4,109.00 $6,170.03 $9,117.00 $13,675.50

OpenStudy (anonymous):

you are right, the equation works for 4,7%

OpenStudy (anonymous):

I believe we're supposed to use the same complicated formula for these problems......

OpenStudy (anonymous):

the one we used for the Evelyn one?

OpenStudy (anonymous):

the one from the 2nd problem. maybe that's why Evelyn result was off

OpenStudy (anonymous):

try to open a new question , maybe someone else is more familiar with deferment payments

OpenStudy (anonymous):

ill see! but thank you so much for your help!

OpenStudy (anonymous):

Eli made a down payment of $1,750 on a jet ski with a purchase price of $9,925. The balance was paid for using a five-year installment loan. If the interest rate on the loan is 11.4%, what is the total cost of the jet ski? $9,106.95 $10,762.80 $12,512.80 $13,066.80 how bout this?